# Contextual Embedding-based Clustering to Identify Topics for Healthcare Service Improvement

### v1

@@ Abstract
Understanding patient feedback is crucial for improving healthcare services, yet analyzing unlabeled short-text feedback presents significant challenges due to limited data and domain-specific nuances. Traditional supervised learning approaches require extensive labeled datasets, making unsupervised methods more viable for uncovering meaningful insights from patient feedback. This study explores unsupervised methods to extract meaningful topics from 439 survey responses collected from a healthcare system in Wisconsin, USA. A keyword-based filtering approach was applied to isolate complaint-related feedback using a domain-specific lexicon. To delve deeper and analyze dominant topics in feedback, we explored traditional topic modeling methods, including Latent Dirichlet Allocation (LDA) and Gibbs Sampling Dirichlet Multinomial Mixture (GSDMM), alongside BERTopic, an advanced neural embedding-based clustering approach. To improve coherence and interpretability where data are scarce and consist of short-texts, we propose kBERT, an integration of BERT embeddings with k-means clustering. Model performance was assessed using coherence scores ($C_{v}$) for topic interpretability and average Inverted Rank-Biased Overlap ($\operatorname{IRBO_{avg}}$) for topic diversity. Results indicate that kBERT achieves the highest coherence ($C_{v}$ = 0.53) and distinct topic separation ($\operatorname{IRBO_{avg}}$ = 1.00), outperforming all other models in short-text healthcare feedback analysis. Our findings emphasize the importance of embedding-based techniques for topic identification and highlight the need for context-aware models in healthcare analytics.
Index Terms:
LDA, GSDMM, BERT, BERTopic, k-means, Clustering, Topic Modeling, Complaint Identification, Healthcare, Short-text
publicationid: pubid: copyright-placeholder
@@ I Introduction
In recent years, online forums have become pivotal in recognizing complaints about services or products. Linguistic studies categorize a complaint as a primary form of communication for expressing dissatisfaction when expectations are not met concerning a product, situation, organization, or event[1]. Both complaints and compliments hold significant weight in business decisions, with complaints being particularly crucial. The analysis of customer product reviews, a practice long utilized in e-commerce, has greatly enhanced business operations and customer satisfaction, ensuring customers feel valued and heard. Similarly, various sectors including banking systems and city utility services are now integrating customer feedback mechanisms to refine their services.
Complaint identification is a nuanced subset of sentiment analysis. Many existing methods have employed supervised machine learning with trained datasets, which are often costly and time-intensive to prepare for domain-specific applications[2]. This challenge has naturally led to an increased interest in unsupervised learning approaches, particularly topic modeling and clustering techniques. These techniques can uncover underlying patterns in textual data without the need for labeled datasets.
@@
Topic modeling aims to identify latent themes within a collection of documents by analyzing word distributions. Methods such as Latent Dirichlet Allocation (LDA) and Gibbs Sampling Dirichlet Multinomial Mixture (GSDMM) have been widely applied to textual data to uncover dominant topics [3]. Clustering methods, in contrast, group similar documents based on their feature representations. Unlike topic models, which allow words to belong to multiple topics, clustering techniques assign each document to a single group based on textual similarity. Recent advances in neural embedding-based clustering, such as BERTopic, have been explored for their potential to handle short and noisy texts compared to LDA and GSDMM. [4, 5]. While topic modeling methods and neural embedding-based clustering approaches have been explored for various text analysis tasks, their effectiveness in scenarios where data is both short and limited remains an open question. In this study, we develop a neural embedding-based clustering model, kBERT, and systematically assess its performance alongside established methods to evaluate their suitability for analyzing short-text on limited data.
This study introduces a system designed to identify topics within patient feedback based on surveys from hospital visits. The system aims to highlight complaint comments by analyzing survey feedback, thereby assisting hospital authorities in enhancing service quality and understanding patient needs more effectively.
Our methodology begins with a manually curated list of keywords to segregate positive and negative comments. The keyword list incorporates domain-specific terms that help capture the nuanced details of patient feedback (see Table II). To complement this keyword-based approach and gain deeper insights into the underlying themes in patient feedback, we employ unsupervised learning methods for topic extraction. We evaluate the performance of traditional topic modeling techniques (LDA, GSDMM) and modern embedding-based clustering approaches (BERTopic) in extracting meaningful topics from a small short-text dataset to better understand the applicability of these methods in patient feedback analysis. Additionally, we compare their results with kBERT to assess its potential in addressing the challenges posed by limited and short healthcare-related textual data. These methods allow for a more comprehensive analysis of patient feedback, addressing both latent topic discovery and document similarity-based clustering to improve topic coherence and interpretability.
Given the challenges associated with short-text length, small dataset sizes, and domain-specific issues in patient feedback, this research aims to explore the following research questions (RQs):
@@
The rest of the paper is structured as follows: Section II reviews related work on complaint identification, topic modeling, and clustering, highlighting the challenges of short-text analysis in healthcare. Section III details the proposed methodology, including data preprocessing, keyword filtering, and applying unsupervised techniques, addressing RQ1. Section IV presents the result analysis, comparing the performance of both topic modeling and clustering approaches, evaluating their coherence, diversity, and interpretability, thereby addressing RQ3. Section V discusses the findings and their implications for healthcare service improvement, effectively answering RQ2. Finally, Section VI concludes the study.
@@ II-A Topic Modeling and Word Embeddings-based Clustering
In the field of NLP, complaint identification, particularly when labeled datasets are available, is often approached as a classification task. This area employs various machine learning models like support vector classifier (SVC), Random Forest, Decision Tree, and multilayer perception (MLP), along with deep learning classifiers such as Dense Neural Networks and convolutional neural networks (CNNs) [6]. Additionally, NLP libraries like VADER, TextBlob, and NLTK are commonly used to derive sentiment scores, enabling the classification of feedback into positive, negative, or neutral categories [7]. However, these libraries might not always capture nuanced details, particularly in domain-specific contexts. With the rise of Large Language Models (LLMs), the field of complaint classification is evolving. L. Yuan et al. introduced a Complaint Text Classification Framework (CCF) based on LLMs, which leverages advanced feature extraction techniques from large language models, significantly enhancing classification accuracy [8].
While classification alone is useful, it doesn’t provide deep insights into the nature of the text. LDA can be effective in understanding the underlying concepts of language. LDA identifies the latent topics within a text corpus. As demonstrated in subsequent studies, LDA is frequently used for sentiment analysis in complaint data. For instance, A. Karami et al. utilized LDA to categorize GEICO customer complaints into four distinct groups [9], while N. Hu et al. applied LDA and Structural Topic Modeling to analyze hotel reviews in New York City, identifying ten major themes of customer dissatisfaction and how these vary across different hotel grades [10]. M. N. Aziz et al. adopted a mixed approach for opinion classification, leveraging both Lexicon-Based (SentiWordNet) and Machine Learning-Based (TF-IDF with SVM) methods, and employed LDA for topic clustering [11]. As research in this field progresses, various topic modeling methods have been compared to evaluate their effectiveness. For instance, H. S. Jung et al. evaluated four topic modeling methods—LDA, Nonnegative Matrix Factorization (NMF), Combined Topic Models (CTM), and BERTopic. Their findings showed that BERT based model demonstrated superior performance across both topic diversity and coherence [12].
Embeddingbased methods have significantly advanced topic modeling by utilizing contextual embeddings like BERT and SBERT to capture semantic and syntactic nuances. The Embedded Topic Model (ETM) [13] integrates LDA with embeddings to enhance topic coherence. Approaches combining BERT embeddings with clustering methods [14, 15] show improved coherence and interpretability. For instance, a framework integrating BERT with clustering and dimensionality reduction [14] demonstrated superior clustering quality, while a hybrid approach combining BERT, LDA [15] achieved high topic coherence. C-Top2Vec [16] further refined topic modeling through multi-vector representations and hierarchical clustering for granular insights.
Topic modeling has evolved with the introduction of advanced models, yet traditional methods still hold significant advantages in certain applications. GPTopic, a novel approach leveraging LLMs, enhances the interpretability and interaction of topic representations, surpassing the traditional “top-word” approach by allowing dynamic modifications and chat-based interactions with topics. However, GPTopic can occasionally generate inaccurate results due to hallucinations, particularly when applied to datasets with fewer than 10,000 documents [17]. While LLMs, including those used in GPTopic, offer flexibility and zero-shot capabilities, they struggle with maintaining topic granularity and often produce overlapping or redundant topics, which complicates the identification of distinct themes [18]. In contrast, traditional methods like LDA and BERT base model remain consistent and reliable, generating coherent topics without hallucinations, making them highly adaptable across various domains without extensive fine-tuning [19].
However, despite their advancements, LLMs require substantial computational resources, making their deployment costly due to the need for high-performance GPUs or TPUs for fine-tuning and inference. This makes traditional methods like LDA and clustering-based approaches more practical for resource-constrained environments. Additionally, LLM-enhanced applications that provide APIs or interfaces like ChatGPT [20], pose risks of data leakage, as models trained on large datasets can unintentionally expose sensitive information. The lack of clarity in how data is stored, processed, and transmitted further exacerbates trust, security, and privacy risks [21, 22].
TABLE I: Datasets
Dataset Name Size Median Type Availability
Patient Feedback 439 14 Short-text private
Clickbait-title 10,000 9 Short-text public
Clickbait-title (500) 500 9 Short-text public
20Newsgroup 15,021 83 Long-text public
20Newsgroup (500) 500 86 Long-text public
@@ II-B Topic Modeling for Short-Text
Conventional methods like LDA encounter challenges due to their reliance on bag-of-words representations, which can be inadequate for capturing the nuances in short texts[23]. Recent research highlights the effectiveness of alternative models. For instance, GSDMM and BERTopic have shown promising results, outperforming LDA in handling short-text. One study by Udupa et al. [23] specifically explored the use of GSDMM and BERTopic for efficient topic clustering in short-text scenarios. Meanwhile, Jiang et al. [24] implemented the Associated Topic Model (ATM) which leverages a collapsed Gibbs sampling algorithm, further emphasizing the diversity of approaches suitable for short-text analysis. Additionally, clustering token-level contextualized word representations from models like BERT has proven effective in capturing polysemy and organizing documents, performing comparably or better than LDA topic models, as found in recent evaluations [25]. These studies collectively suggest a shift towards models that can better interpret the limited contextual information available in shorter documents.
@@ II-C Topic Modeling in Healthcare Feedback Service
TABLE II: Sample Keyword (Hot Word) List
infect angry malpractice delay MRSA
blood sepsis cops grievance mis
trauma wound abus argu un
negl viol massac hipaa hippa
african american gun bruise burn
@@ III Methodology
This study focuses on short-text patient feedback, specifically in small datasets, where traditional topic modeling techniques often face challenges. We apply keyword filtering to identify complaint-related feedback using a curated lexicon, ensuring a focused analysis on patient grievances. Simultaneously, topic modeling is employed to extract dominant themes, providing insights into both complaints and general concerns.
As our dataset consists of short-text feedback, GSDMM is used as the baseline for short-text topic modeling, while LDA is included as a reference for long-text topic modeling. These models are compared with BERTopic and kBERT, which leverage transformer-based embeddings to enhance topic coherence in small short-text datasets.
@@ III-A Dataset Description
To evaluate the robustness of our methodology, we also tested it on two additional public datasets: the Clickbait-title dataset [29] and the 20Newsgroup dataset[30]. The Clickbait-title dataset is a short-text dataset comprising headlines categorized as clickbait or non-clickbait. We selected 10,000 samples from this dataset, with a median word count of 9, making it representative of short-text content. Additionally, to examine model performance on small short-text datasets, we extracted a subset of 500 samples (clickbait-title (500)). In contrast, the 20Newsgroup dataset consists of longer articles from 20 distinct topics. It is a widely used benchmark dataset for different natural language tasks. To standardize the data, we excluded the bottom and top 10% of articles based on word count and retained 15,021 articles for evaluation and the median word count is 83. We also extracted a subset of 500 samples (20Newsgroup (500)) to explore small dataset scenarios in long-text settings.
These datasets allow for a comprehensive evaluation of topic modeling and clustering techniques across varying text lengths (short vs. long) and dataset sizes (large vs. small), providing insights into how different approaches perform under diverse conditions.
@@ III-B Keyword Processing
In this work, we utilize a carefully curated keyword list to effectively filter and categorize the patient feedback into positive and negative comments. These keywords, which are crucial for analyzing the sentiment of the feedback, are presented in Table II. This lexicon ensures a comprehensive reflection of the terminology commonly found in patient complaints, aiding in the accurate assessment of their sentiment. The keyword list is also important as it contains domain-specific nuances. Some feedback includes medical and legal terms that may not explicitly mention complaints. For instance, “I saw another patient’s information on the clinic’s screen, which seems like a HIPAA issue”. These subtle details are often missed by general-purpose models. Capturing these nuances is crucial for improving patient care. In Table II, the keywords are systematically classified into four distinct groups to enhance the precision of complaint detection in the texts. This classification is as follows:
  • Full Keywords: These encompass complete terms that are often used in their entirety within complaints, such as “accident”, “delay”, “agitate”, “american” etc.
  • Shortened Keywords: Truncated forms of words, like “abus” (for abuse) and “argu” (for argue), which are used in different parts of speech (POS) in the feedback.
  • Prefix Keywords: Terms like “mis” and “un”, serve as prefixes, indicative of negative connotations when attached to other words.
  • Jargon/Medical Term Keywords: This category includes essential terms like “sepsis”, “bruise”, “HIPAA”, and “MRSA”, crucial for identifying comments on various topics such as regulatory compliance or infection-related concerns. ‘HIPPA’ is included as it is a frequent misspelling of ‘HIPAA’ by patients.
Each group is designed to address the different ways keywords could manifest in the data — as standalone words, abbreviated forms, or as part of medical jargon — thus providing a comprehensive filter for the preliminary categorization of comments.
In our methodology, the keyword processing phase presented a unique challenge — the ambiguity arising from shortened keywords. For example, the abbreviation ‘compl’ is extracted from terms such as ‘complaint’ and ‘complete’, though only the former is relevant to our analysis. To mitigate this, we refined our keyword list to include only the full, unambiguous forms of each word, ensuring that only contextually relevant terms are considered.
Another complexity arose from the morphological forms of keywords within the dataset. Words like ‘argue’ appeared in various derivatives, such as ‘argues’, ‘argued’, or ‘arguing’. Traditional lemmatization is employed to address this, which involves reducing words to their base or dictionary form. Lemmatization’s effectiveness is contingent upon accurate POS tagging. The process needs to recognize the grammatical role of words within the context of their use. This is because POS tagger models are typically trained on complete sentences or documents rather than isolated tokens. This training approach poses a challenge when a word like ‘argument’ needs to be reduced to ‘argue’ [31].
To navigate these linguistic nuances, we utilize the ‘word_forms’ library in Python [32]. This library provides a comprehensive mapping of word forms across different POS categories: nouns, verbs, adverbs, and adjectives; allowing us to accurately identify and process the various permutations of a keyword in our dataset. The complete keyword processing procedure is detailed in Algorithm 1.
@@ III-C Feedback Preprocessing
The fidelity of topic modeling in NLP is heavily reliant on the quality of input data. Recognizing this, our methodology is built on a thorough protocol for preprocessing data. Initially, tokenization—the segmentation of text into individual terms—is conducted to granulate the dataset into analyzable units.
Subsequent to tokenization, we focus on the elimination of stop words. These words, while structurally integral, contribute minimally to the semantic load of the text. Common stop words such as ‘a’, ‘and’, ‘the’, and ‘of’ are identified and removed. Their high frequency and low informational value make them prime candidates for exclusion to enhance data purity. To perform this task, we employ the ‘spacy’ library—a powerful and versatile Python toolkit for advanced NLP [33]. Its sophisticated algorithms allowed for an efficient and thorough filtration process, excising stop words from the tokenized text. This cleansing step is vital, as it significantly reduces noise within the data, thereby streamlining the path for more accurate and meaningful topic modeling. The detailed procedure for this filtration is presented in Algorithm 2.
@@ III-D1 Latent Dirichlet Allocation (LDA)
LDA is a generative probabilistic model widely used for uncovering latent topics within a collection of text documents [34]. LDA assumes that each document is a mixture of various topics, where each topic is characterized by a probability distribution over words. By breaking down a document-word matrix into two smaller matrices—the Document-Topic Matrix and the Topic-Word Matrix—LDA reveals the relationships between documents and their underlying topics, as well as the association between topics and the defining words. This process makes LDA particularly valuable for analyzing large collections of text data. The LDA model operates through a hierarchical generative process. At the document level, a topic distribution $\theta$ for each document $d$ is drawn from a Dirichlet distribution with parameter $\alpha$, which serves as the prior for the document-topic distributions. For each word $w_{d,n}$ in a document, a topic $z_{d,n}$ is sampled from the document’s topic distribution $\theta_{d}$. Subsequently, a word is generated by sampling from the topic-specific word distribution $\beta_{z_{d,n}}$, which is drawn from another Dirichlet distribution with parameter $\eta$. This layered structure captures both the document-level topic diversity and the word-level thematic coherence.
Figure 1: LDA Blueprint (Source:[13, 35] )
In our study, we applied LDA to analyze hospital patient feedback documents, identifying and categorizing latent themes within the text. Each document was modeled as a probabilistic mixture of topics, with $\theta_{d,k}$ representing the likelihood of topic $k$ in document $d$, and $\beta_{k,w}$ indicating the relevance of word $w$ within topic $k$. To enhance thematic clarity, we experimented with 3 to 20 topic clusters to balance specificity and breadth. The Dirichlet parameter $\alpha$ was set to ‘auto,’ allowing dynamic adjustment of document-topic distributions. We also fine-tuned the top words per topic, improving the model’s ability to highlight relevant terms and themes.

### v3

@@ Abstract
Understanding patient feedback is crucial for improving healthcare services, yet analyzing unlabeled short-text feedback presents challenges due to limited data and domain-specific nuances. Traditional supervised approaches require extensive labeled datasets, making unsupervised methods more practical for extracting insights. This study applies unsupervised techniques to analyze 439 survey responses from a healthcare system in Wisconsin, USA. A keyword-based filter was used to isolate complaint-related feedback using a domain-specific lexicon. To identify dominant themes, we evaluated traditional topic models such as Latent Dirichlet Allocation (LDA) and Gibbs Sampling Dirichlet Multinomial Mixture (GSDMM) - alongside BERTopic, a neural embedding-based clustering method. To improve coherence and interpretability in sparse, short-text data, we propose kBERT, which integrates BERT embeddings with k-means clustering. Model performance was assessed using coherence scores ($C_{v}$) and average Inverted Rank-Biased Overlap ($\operatorname{IRBO_{avg}}$). kBERT achieved the highest coherence ($C_{v}$ = 0.53) and topic separation ($\operatorname{IRBO_{avg}}$ = 1.00), outperforming all other models. These findings highlight the value of embedding-based, context-aware models in healthcare analytics.
@@ I Introduction
Understanding and addressing complaints is essential across domains, including healthcare. Linguistic studies define a complaint as a form of communication expressing dissatisfaction when expectations are unmet [17]. While industries like e-commerce and utilities increasingly use feedback to improve services, healthcare systems are also beginning to leverage patient comments for quality enhancement. Complaint identification is often approached through supervised learning, but such methods require large labeled datasets, which are resource-intensive to build. As a result, unsupervised techniques like topic modeling and clustering have gained attention for their ability to extract meaningful insights from unlabeled text.
@@
Topic modeling aims to identify latent themes within a collection of documents by analyzing word distributions. Methods such as Latent Dirichlet Allocation (LDA) and Gibbs Sampling Dirichlet Multinomial Mixture (GSDMM) have been widely applied to textual data to uncover dominant topics [7]. Clustering methods, in contrast, group similar documents based on their feature representations. Unlike topic models, which allow words to belong to multiple topics, clustering techniques assign each document to a single group based on textual similarity. Recent advances in neural embedding-based clustering, such as BERTopic, have been explored for their potential to handle short and noisy texts compared to LDA and GSDMM. [8]. While topic modeling methods and neural embedding-based clustering approaches have been explored for various text analysis tasks, their effectiveness in scenarios where data is both short and limited remains an open question. In this study, we develop a neural embedding-based clustering model, kBERT, and systematically assess its performance alongside established methods to evaluate their suitability for analyzing short-text on limited data.
This study presents a system for analyzing patient survey feedback from hospital visits, with an emphasis on identifying complaint-related topics. To address the challenges of short, sparse, and domain-specific text, we use a curated keyword list to filter negative feedback and apply unsupervised methods to uncover dominant themes. We evaluate traditional topic models (LDA, GSDMM) and embedding-based approaches (BERTopic, kBERT) on a small short-text dataset. By combining domain-aware filtering with neural embeddings, our approach aims to improve topic coherence and interpretability in low-resource healthcare settings. This work is guided by the following research questions (RQs):
@@ II-A Topic Modeling and Word Embeddings-based Clustering
In NLP, complaint identification is often treated as a classification task when labeled data is available. Traditional approaches use machine learning models such as SVC, Random Forest, and MLP, as well as deep learning methods like CNNs. While basic sentiment analysis tools (e.g., VADER, TextBlob, NLTK) can classify feedback as positive, negative, or neutral, they often miss domain-specific nuances. Recent advancements in large language models (LLMs) have significantly improved various NLP tasks [16]. Frameworks like the Complaint Text Classification Framework (CCF) leverage LLM-based feature extraction to enhance classification accuracy [23]. While classification is helpful, it often lacks deeper interpretability. Topic modeling techniques like LDA are widely used to uncover latent themes in textual data and have been applied across domains to analyze customer feedback. As the field evolves, several studies have compared topic modeling methods to assess their effectiveness. For example, Jung et al. evaluated LDA, NMF, CTM, and BERTopic, finding that BERT-based models outperformed others in both topic diversity and coherence [12]. Embedding-based methods have significantly advanced topic modeling by utilizing contextual embeddings like BERT and SBERT to capture semantic and syntactic nuances. The Embedded Topic Model (ETM) [9] integrates LDA with embeddings to enhance topic coherence.
@@ II-B Topic Modeling for Short-Text
Conventional methods like LDA struggle with short texts due to their reliance on bag-of-words representations, which often fail to capture contextual nuances [21]. Recent studies have shown that alternative models like GSDMM and BERTopic outperform LDA in short-text scenarios. For example, Udupa et al. demonstrated the effectiveness of both GSDMM and BERTopic for topic clustering in brief texts [21]. Additionally, clustering token-level contextualized embeddings from models like BERT has proven effective in capturing polysemy and organizing documents, often matching or exceeding LDA in performance [20]. These findings indicate a shift toward models better suited for interpreting limited-context data.
@@ II-C Topic Modeling in Healthcare Feedback Service
TABLE I: Sample Keyword (Hot Word) List
Type Examples
Full angry malpractice delay grievance african
Short negl abus argu negl Viol
Jargon hipaa hippa MRSA
Prefix un mis
@@ III Methodology
This study analyzes short-text patient feedback using keyword filtering and unsupervised topic modeling. GSDMM serves as the baseline for short-text modeling, while LDA provides a long-text reference. We also implement BERTopic and kBERT, which leverage transformer-based embeddings to enhance coherence.
@@ III-A Dataset Description
To evaluate the robustness of our methodology, we also tested it on two public datasets: the Clickbait-title dataset [6] and the 20Newsgroup dataset [1]. The Clickbait-title dataset contains short headlines labeled as clickbait or non-clickbait. We selected 10,000 samples with a median word count of 9, and also created a smaller subset of 500 samples to evaluate performance on small short-text data. The 20Newsgroup dataset consists of longer articles across 20 distinct topics and is widely used for benchmarking NLP tasks. A 500-sample subset was extracted to simulate small long-text scenarios. These datasets allow for a comprehensive evaluation of topic modeling and clustering techniques across varying text lengths (short vs. long) and dataset sizes (large vs. small), providing insights into how different approaches perform under diverse conditions.
@@ III-B Keyword Processing
In this work, we utilize a carefully curated list of domain-specific keywords to filter patient feedback into complaint and non-complaint categories. These keywords capture subtle indicators of dissatisfaction, such as legal, medical, or emotional terms that are often overlooked by general-purpose models. We grouped the keywords into four categories: full, shortened, prefix, and jargon. These categories, along with representative examples, are shown in Table I. For instance, a comment like “I saw another patient’s information on the clinic’s screen, which seems like a HIPAA issue” highlights a potential privacy concern without explicitly using negative sentiment words, which domain-specific filtering can effectively capture.
A key challenge in keyword processing is handling ambiguous terms, especially shortened forms. For example, the abbreviation compl could represent either complaint or complete, but only the former is relevant in our context. To avoid such ambiguity, we refined the list to include only unambiguous and contextually meaningful terms. Another complexity arises from morphological variations of keywords. Words like argue may appear as argues, argued, or arguing in the dataset. Addressing this requires effective lemmatization, which reduces words to their base forms. However, the success of lemmatization depends heavily on accurate part-of-speech (POS) tagging. Most POS taggers are trained on full sentences rather than isolated tokens, making context-sensitive lemmatization particularly challenging [3].
To manage these linguistic nuances, we employ the word_forms Python library. This tool provides comprehensive mappings of words across POS categories—nouns, verbs, adverbs, and adjectives—allowing us to robustly identify and process all morphological variants of keywords in our dataset. The complete keyword processing procedure is detailed in Algorithm 1.
@@ III-C Feedback Preprocessing
The effectiveness of topic modeling in NLP depends heavily on the quality of input data. To ensure this, our methodology follows a structured preprocessing pipeline. We begin with tokenization, which segments text into individual terms to make it analyzable. Next, we remove stop words such as “a,” “and,” “the,” and “of,” which are frequent but carry little semantic value. This step helps reduce noise and improve data clarity. We use the “spacy” library for stop word removal due to its robust and efficient NLP capabilities. This preprocessing step is essential for producing cleaner input that enhances the accuracy and interpretability of downstream topic modeling. The detailed steps are outlined in Algorithm 2.
@@ III-D1 Latent Dirichlet Allocation (LDA)
LDA is a generative probabilistic model that uncovers latent topics within a document collection by modeling each document as a mixture of topics, where each topic is a distribution over words [4]. It decomposes the document-word matrix into a document-topic and a topic-word matrix, revealing underlying thematic structures. At the core of LDA, topic proportions for each document are drawn from a Dirichlet prior, and words are generated based on topic assignments and word distributions. In our study, we applied LDA to hospital patient feedback to identify latent themes. We experimented with 3 to 20 topic clusters to balance specificity and generality, using the auto setting for Dirichlet priors to adapt topic distributions. Top words per topic were fine-tuned to enhance interpretability.